# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws (Extended Version)

### v2

@@ Acknowledgment
Appendix A An Example of Application of the Proposed State-Centric Necessary Condition
@@
Let’s introduce Corollary 4 of order $n$ as follows. We will show why it is challenging to derive the existence/non-existence of chattering based on PMP-based necessary conditions, and how to obtain the result based on the proposed state-centric necessary condition. Specifically, we will try to derive the result based on PMP-based/state-centric necessary conditions as examples.
A-A Problem Statement
@@ A-B Preliminaries
The proof of the above conclusions can be found in our related work [32] (Section IV-B).
@@
A-C Why Deriving the Existence/Non-Existence of Chattering through PMP-Based Necessary Conditions is Challenging?
@@
$\begin{dcases}\dot{\lambda}_{2}=-\lambda_{3}+\eta,\\ \dot{\lambda}_{n}=0,\\ \dot{\lambda}_{k}=-\lambda_{k+1},\,k\in\left\{1,3,4,5,\dots,n-1\right\}.\\ \end{dcases}$ (67)
@@
$u\left(t\right)=\begin{dcases}-1,&\lambda_{1}\left(t\right)>0,\\ 0,&\lambda_{1}\left(t\right)\equiv 0\text{ holds for a period},\\ 1,&\lambda_{1}\left(t\right)<0.\\ \end{dcases}$ (69)
@@
where $p_{i}$ is a polynomial of order no higher than $n-3$. Furthermore,
@@
$\mathcal{C}=\left\{\begin{aligned} &\left(\left(\mu_{i}\right)_{i\in{\mathbb{N}}^{*}},\left(t_{i}\right)_{i\in{\mathbb{N}}},\left(p_{i}\left(\bullet\right)\right)_{i\in{\mathbb{N}}},\left(\lambda_{1i}\left(\bullet\right)\right)_{i\in{\mathbb{N}}}\right):\\ &\qquad t_{i}<t_{i+1},\,\lim_{i\to\infty}t_{i}<\infty,\,\lambda_{0n}\not=0,\,\mu_{i}\geq 0,\\ &\qquad p_{i}\left(\bullet\right)\text{ is a polynomial of order no higher than }n-3,\\ &\qquad\forall t\in\left(t_{i-1},t_{i}\right),\,p_{i}\left(t\right)>0,\\ &\qquad\forall t\in{\mathbb{R}},\,\lambda_{1i}\left(t\right)=\lambda_{0n}\frac{\left(-1\right)^{n-1}}{\left(n-1\right)!}\left(t-\frac{1}{4}t_{i}-\frac{3}{4}t_{i-1}\right)\left(t-\frac{3}{4}t_{i}-\frac{1}{4}t_{i-1}\right)p_{i}\left(t\right),\\ &\qquad\forall t\in{\mathbb{R}},\,\lambda_{1\,i+1}\left(t\right)=\lambda_{1i}\left(t\right)+\mu_{i}\left(t-t_{i}\right)\end{aligned}\right\}.$ (79)
@@
A-D Why Deriving the Existence/Non-Existence of Chattering through the Proposed State-Centric Necessary Condition is Possible?
@@
Once the initial value $\left(\tau_{i}\right)_{i=1}^{n-2}$ is given, the whole sequence $\left(\tau_{i}\right)_{i\in\infty}$ is determined. In other words, a sufficient condition for the non-existence of chattering is $\widehat{\mathcal{C}}=\varnothing$, where
@@
A-E Deriving the Non-Existence of Chattering through the Proposed State-Centric Necessary Condition in 4th-Order Problem
Based on the proposed state-centric necessary condition, we can even directly prove that $\widehat{\mathcal{C}}=\varnothing$ when $n=4$; hence, chattering does not exist in the optimal control of problem (62). Note that the above claim is a novel theoretical result.
@@
$=$ $144\left(\tau_{i-1}-\tau_{i-3}\right)\left(\tau_{i}-\tau_{i-2}\right)f\left(\tau_{i-1}-\tau_{i};\tau_{i-2}-\tau_{i-1},\tau_{i-3}-\tau_{i-2}\right)=0,$
@@
(88) has two roots in $\left(0,r_{i}\right)$ and $\left(1,+\infty\right)$, respectively. Hence, $10<r_{i+1}<r_{i}<$ 1. Denote $r_{\infty}\triangleq\lim_{i\to\infty}r_{i}\in\left[0,1\right]$. Then, $g\left(r_{\infty},r_{\infty}\right)$ implies that $r_{\infty}=0$. Note that

### v3

@@ A-D Proofs in Section VI
Appendix B An Example of Application of the Proposed State-Centric Necessary Condition
@@
Let’s introduce Corollary 4 of order $n$ as follows. We will show why it is challenging to derive the existence/non-existence of chattering based on PMP-based necessary conditions, and how to obtain the result based on the proposed state-centric necessary condition. Specifically, we will try to derive the result based on PMP-based/state-centric necessary conditions as examples. Note that the fundamental idea of the state-centric approach has already been applied in our previous work [31, Section VI.B].
B-A Problem Statement
@@ B-B Preliminaries
The proof of the above conclusions can be found in our related work [31, Section VI.B].
@@
B-C Why Deriving the Existence/Non-Existence of Chattering through PMP-Based Necessary Conditions is Challenging?
@@
[…]
@@
[…]
@@
where $p_{i}$ is a polynomial of order no higher than n3n-3. Furthermore,
@@
[…]
@@
B-D Why Deriving the Existence/Non-Existence of Chattering through the Proposed State-Centric Necessary Condition is Possible?
@@
Once the initial value $\left(\tau_{i}\right)_{i=1}^{n-2}$ is given, the whole sequence $\left(\tau_{i}\right)_{i\in\infty}$ is determined. In other words, a sufficient condition for the non-existence of chattering is $\widehat{\mathcal{C}}$$\varnothing$, where
@@
B-E Deriving the Non-Existence of Chattering through the Proposed State-Centric Necessary Condition in 4th-Order Problem
Based on the proposed state-centric necessary condition, we can even directly prove that $\widehat{\mathcal{C}}$$\varnothing$ when $n=4$; hence, chattering does not exist in the optimal control of problem (52). Note that the above claim is a novel theoretical result.
@@
[…]
@@
(78) has two roots in $\left(0,r_{i}\right)$ and $\left(1,+\infty\right)$, respectively. Hence, $0<r_{i+1}<r_{i}<1$. Denote $r_{\infty}\triangleq\lim_{i\to\infty}r_{i}\in\left[0,1\right]$. Then, $g\left(r_{\infty},r_{\infty}\right)$ implies that $r_{\infty}=0$. Note that